# Caging in Time: A Framework for Robust Object Manipulation under Uncertainties and Limited Robot Perception

### v3

@@ 7 Experiments
To quantify manipulation precision, we adopted the Mean Absolute Error (MAE) as our primary metric. This was calculated by averaging the absolute distances between the actual and reference trajectories throughout the manipulation process. Through these experiments, we aimed to thoroughly assess the performance, robustness, and versatility of our Caging in Time framework across a range of quasi-static and dynamic manipulation tasks.
@@ 7.1 Quasi-static Tasks
Through quasi-static experiments, we would like to investigate the performance of our framework from three aspects: 1) Without any sensing feedback besides the size and initial position of the bounding circle, can robust planar manipulation be achieved by the proposed Caging in Time? 2) If so, what precision can be achieved, and how is the robustness affected by different settings of the framework and outer disturbances? 3) Under imperfect perceptions like positional noise and network lag, can Caging in Time surpass closed-loop methods?
As displayed in Fig. 14, The robot in this task used a 100 mm line pusher, and objects of various shapes were 3D-printed in a similar size, fitting within a bounding circle of radius $r=25$ mm. The pushing distance of each action was set as $d_{push}=20$ mm.
@@ 7.1.1 Evaluation of Cage Settings
Figure 17: Robustness validation with a $\infty$-shaped trajectory. This experiment was conducted with the heart-shaped object (9), recorded by camera 2. Top: Ten loops conducted by the Caging in Time framework. Bottom: A failed attempt by a naïve pushing strategy.
@@ 7.1.2 Why Caging in Time
7.1.3 Comparison with a Closed-loop Method
We introduced two types of disturbances to simulate real-world scenarios with imperfect perception and communication delays. Firstly, positional noise was added as Gaussian noise with a standard deviation of the noise level in both $x$ and $y$ directions, simulating the effect of imperfect perception, such as that from an RGBD camera under dim lighting conditions. Secondly, to emulate potential packet loss during data transmission or sudden obstruction of the object from the camera’s view, we introduced a network lag that for each execution step, there was a $50\%$ probability that the controller would receive the object’s position from 0.5 or 1 second ago.
@@ 7.2 Dynamic Tasks
In the following experiments, the plate followed a predefined translational trajectory $\mathcal{T}=\{\mathbf{x}_{p}(t)\}_{t=0}^{T}$. While the trajectory of the plate was preset, the tilt angle of the plate at each time step is computed by our framework. A Cartesian space motion control is then used to control the robot’s end-effector to follow this trajectory in a pure open-loop manner. For all dynamic experiments, we also used MAE to quantify the performance. Since the plate accurately tracks the target trajectory, the MAE is calculated as the average distance between the ball and the center of the plate across all time steps and all trials.
@@ 7.2.2 Ball Catching
In the second experiment, we introduced greater uncertainty by having a human manually toss the ball onto the plate, as illustrated in Fig. 22 and Fig. 2. Due to the inherent difficulty in controlling the landing position of hand-tossed balls, we only recorded trials where the ball successfully landed on the plate. Out of 20 such successful tosses, the system was able to catch and stabilize the ball in 13 trials. This partial success rate demonstrates the framework’s robustness to velocity uncertainties within a certain range. The partial success rate in catching human-tossed balls demonstrates the limitations of the Caging in Time framework’s pre-configured error tolerance. This tolerance is constrained to ensure feasible actions within the robot’s hardware torque limits, and some human tosses evidently introduced velocity variations beyond this tolerance range.
@@ 7.2.3 Caging with Uncertainty
To note that, The success of these tasks can be partially attributed to the nature of irregular shapes, which are less likely to roll freely due to their inherent energy traps created by their non-uniform geometry. This characteristic actually aids in maintaining stability, as the objects tend to settle into local energy minima, complementing the caging strategy of our framework.
@@ 7.2.4 Caging in Higher Dimensions
Figure 24: Ball balancing using Caging in Time with the dimension $n=2$ while tracing “RICE”. Left: Representative frames from video recordings. The overlaid transparent balls show the ball’s position at different time points during the task. Right: Qualitative trajectory plots of 10 repeated trials for each task.
@@ 8 Conclusion
In this work, we proposed and evaluated Caging in Time, a novel theory for robust object manipulation. Our framework demonstrated high precision in both quasi-static and dynamic tasks without requiring detailed object information or real-time feedback. Rigorous evaluations highlighted the framework’s resilience and adaptability to various objects and dynamic scenarios. The Caging in Time approach proved effective in handling new objects, positional perturbations, and challenging dynamic tasks, showcasing its potential for reliable manipulation in uncertain environments.
@@
Looking forward, we aim to address these limitations and expand the horizons of Caging in Time. A key direction for future research is the integration of learning techniques and Large Language Models (LLMs) to autonomously discover and learn Caging in Time tasks from everyday scenarios,which could enable the framework to automatically derive appropriate propagation functions and strategies. We also plan to explore its application to multi-object scenarios and more complex cases, potentially leveraging machine learning to handle increased task complexity and environmental variability.
As the Caging in Time theory develops with more generalizable algorithms to handle increasingly complex task dynamics, we anticipate enabling many interesting and robust real-world applications. These may include in-hand manipulation with compliant robot hands, robot picking from clutter by pushing against environmental constraints, and dynamic transportation of large objects in robot hands without secured grasps. The framework’s demonstrated ability to handle both quasi-static and dynamic tasks opens up new possibilities for robust, open-loop control strategies in various robotic manipulation scenarios.
@@ 8.1.1 The Peshkin Distance
where $a$ represents the radius of the object’s circumscribing circle centered at the mass center, $c$ is the distance from the mass center to the contact point, and $\beta$ is the angle subtended by the motion direction and the line connecting the contact point with the mass center. The angles $\beta_{0}$ and $\beta_{1}$ correspond to the initial and final orientation angles for this rotation.
@@
$=\frac{mc^{2}}{a^{2}+c^{2}}\sin{\beta_{0}}\cos{\beta_{0}}$
$\leq m\sin{\beta_{0}}\cos{\beta_{0}}\leq\frac{m}{2}$
@@
$\mathcal{V}_{q_{t}}=U(q_{t},u_{t})=$ (32)
if $\mathbf{d}(q_{t},\mathcal{L})>r+d_{push}$ or $u_{t}$ is None:
@@ 8.2.1 System Dynamics
The system dynamics of the ball on the tilting plate can be described by the following equation:
@@
$=\frac{m(1+\eta_{m})}{m+\frac{I_{b}}{r_{b}^{2}}}\Big{(}(I_{n+1}-\mathbf{n}(\boldsymbol{\theta})\mathbf{n}(\boldsymbol{\theta})^{T})\times(\mathbf{g}_{\theta}+\mathbf{a}_{p}+\boldsymbol{\eta}_{p})\Big{)}-(\mu_{r}+\eta_{\mu})\dot{\mathbf{x}}_{t}$
@@
$f_{t}=$ $\frac{m}{m+\frac{I_{b}}{r_{b}^{2}}}(1+\eta_{m})((g+(\ddot{z}_{p}+\eta_{z}))\sin\theta+(\ddot{x}_{p}+\eta_{x})\cos\theta)-(\mu_{r}+\eta_{\mu})\dot{x}_{t}$ (37)
$g_{t}=$ $\frac{m}{m+\frac{I_{b}}{r_{b}^{2}}}(1+\eta_{m})((g+(\ddot{z}_{p}+\eta_{z}))\cos\theta-(\ddot{x}_{p}+\eta_{x})\sin\theta)$

### v4

@@ 7 Experiments
To quantify manipulation precision, we adopted the Mean Absolute Error (MAE) as our primary metric, calculated by averaging the absolute distances between the actual and reference trajectories throughout the manipulation process.
As this initial exposition of our framework, we deliberately selected two representative manipulation paradigms — one quasi-static and one dynamic — to evaluate fundamental capabilities while facilitating clear visualization of PSS propagation. Through these experiments, we aimed to thoroughly assess the performance, robustness, and versatility of our Caging in Time framework, establishing a foundation for future extensions to more scenarios.
@@ 7.1 Quasi-static Tasks
Through quasi-static experiments, we aimed to investigate the performance of our framework from three aspects: 1) Without any sensing feedback besides the size and initial position of the bounding circle, can robust planar manipulation be achieved by the proposed Caging in Time? 2) If so, what precision can be achieved, and how is the robustness affected by different settings of the framework and outer disturbances? 3) Under imperfect perceptions like positional noise and network lag, can Caging in Time surpass closed-loop methods?
As displayed in Fig. 14, the robot in this task used a 100 mm line pusher, and objects of various shapes were 3D-printed in a similar size, fitting within a bounding circle of radius $r=25$ mm. The pushing distance of each action was set as $d_{push}=20$ mm. The computation time for each step in action selection and PSS propagation was $26.2\pm 8.5$ ms.
@@ 7.1.1 Evaluation of Cage Settings
Figure 17: Robustness validation with a $\infty$-shaped trajectory. This experiment was conducted with the heart-shaped object (9), recorded by camera 2. Top: Ten loops conducted by Caging in Time. Bottom: A failed attempt by a naïve pushing strategy.
@@ 7.1.2 Why Caging in Time
7.1.3 Comparison with a Closed-loop Method
We introduced two types of disturbances to simulate real-world scenarios with imperfect perception and communication delays. Firstly, positional noise was added as Gaussian noise with a standard deviation of the noise level in both $x$ and $y$ directions, simulating the imperfect perception, such as that from an RGBD camera under dim lighting conditions. Secondly, to emulate potential packet loss during data transmission or sudden obstruction of the object from the camera, we introduced a network lag that for each execution step, there was a $50\%$ possibility that the controller would receive the object’s position from 0.5 or 1 second ago.
@@ 7.2 Dynamic Tasks
Figure 21: Ball catching success rate evaluation of finding feasible action sequences in PSS propagation over initial velocity uncertainty range $\Delta\dot{\mathbf{x}}_{0}$, average initial ball velocity $\overline{\dot{\mathbf{x}}}_{0}$, and hardware limitation represented by the end-effector angular acceleration limit $\beta_{max}$.
@@
In the following experiments, the plate followed a predefined translational trajectory $\mathcal{T}=\{\mathbf{x}_{p}(t)\}_{t=0}^{T}$. While the trajectory of the plate was preset, the tilt angle of the plate at each time step is computed by our framework. A Cartesian space motion control is then used to control the robot’s end-effector to follow this trajectory in a pure open-loop manner. For all dynamic experiments, we also used MAE to quantify the performance. Since the plate accurately tracks the target trajectory, the MAE is calculated as the average distance between the ball and the center of the plate across all time steps and all trials. The computation time for each step in action selection and PSS propagation was $67.3\pm 12.6$ ms.
7.2.1 Dynamic Sensitivity Analysis
Before proceeding with experiments, we conducted a dynamic sensitivity analysis to examine how Caging in Time performance in the ball catching tasks responds to initial state distribution and hardware limitations, as previously mentioned in Alg. 1 and Eq. 26. Fig. 21 illustrates the success rates for finding feasible action sequences in Alg. 1 for a ball catching task across varying initial velocity uncertainties $\Delta\dot{\mathbf{x}}_{0}$, average initial velocities $\overline{\dot{\mathbf{x}}}_{0}$, and end-effector angular acceleration limits $\beta_{max}$. The left plot fixes $\overline{\dot{\mathbf{x}}}_{0}$ at 0.8m/s, demonstrating how acceleration capabilities counteract uncertainty effects, while the right plot fixes $\Delta\dot{\mathbf{x}}_{0}$ at 0.05m/s, revealing velocity-hardware requirement relationships. We sampled each parameter using 10 evenly-spaced values within their ranges, with 100 trials per configuration done in simulated PSS propagation in Alg. 1.
As shown in Fig. 21, with a fixed $\overline{\dot{\mathbf{x}}}_{0}$ of 0.8 m/s, the tolerance of $\Delta\dot{\mathbf{x}}_{0}$ improved to 0.2 m/s as $\beta_{max}$ increased beyond 20 rad/$\text{s}^{2}$, indicating that higher acceleration capabilities allow for the system to accommodate larger PSS. Similarly, with $\Delta\dot{\mathbf{x}}_{0}$ fixed at 0.05 m/s, higher $\overline{\dot{\mathbf{x}}}_{0}$ required proportionally higher $\beta_{max}$ to maintain 100% success rates. Based on our Franka robot’s maximum end-effector angular acceleration, corresponding to the robot configuration used for this task, of approximately 25 rad/$\text{s}^{2}$, we determined our setup could reliably handle initial velocities up to 0.9 m/s with uncertainty ranges below 0.15 m/s, which guided our subsequent experimental design for dynamic tasks.
@@ 7.2.3 Ball Catching
In the second experiment, we introduced greater uncertainty by having a human manually toss the ball onto the plate, as illustrated in Fig. 23 and Fig. 1. Due to the inherent difficulty in controlling the landing position of hand-tossed balls, we only recorded trials where the ball successfully landed on the plate. Out of 20 such successful tosses, the system was able to catch and stabilize the ball in 13 trials. This partial success rate demonstrates the framework’s robustness to velocity uncertainties within a certain range, as well as the limitations of the Caging in Time framework’s pre-configured error tolerance. This tolerance is constrained to ensure feasible actions within the robot’s hardware torque limits, and some human tosses evidently introduced velocity variations beyond this tolerance range.
To quantitatively evaluate task completion rates under varying uncertainty conditions, we analyzed success rates in ball catching across different initial velocity distributions. Our data shows that slope dropping with velocities of $0.84\pm 0.12$ m/s achieved a 100% success rate, while human tossing with higher variability beyond the tolerance range ($0.93\pm 0.28$ m/s) yielded a 65% success rate. This real-world evaluation further validates our dynamic sensitivity analysis in Fig. 21 and demonstrate the framework’s performance boundaries under different uncertainty magnitudes.
@@ 7.2.4 Caging with Uncertainty
To note that, The success of these tasks can be partially attributed to the nature of irregular shapes, which are less likely to roll freely due to their inherent energy traps created by their non-uniform geometries. This characteristic actually aids in maintaining stability, as the objects tend to settle into local energy minima, complementing the caging strategy of our framework.
@@ 7.2.5 Caging in Higher Dimensions
Figure 25: Ball balancing using Caging in Time with the dimension $n=2$ while tracing “RICE”. Left: Representative frames from recordings. The overlaid transparent balls show the ball’s position at different time points during the task. Right: Qualitative trajectory plots of 10 repeated trials for each task.
8 Applications in practice
Applications for Caging in Time could potentially extend beyond previous experiments to broader manipulation domains. This section explores the possibilities and practical considerations for implementing Caging in Time in various scenarios.
8.1 Example Applications
In Caging in Time, according to Sec. 3.2, object states are represented through PSS $\mathcal{Q}_{t}$ within task-specific state spaces, with actions $u_{t}$ ensuring $\mathcal{Q}_{t+1}=\Pi(\mathcal{QV}_{t})\subset\mathcal{S}^{t+1}_{cage}$, where $\mathcal{QV}_{t}$ is the motion bundle in Eq. (4) and $\Pi$ is the propagation function in Eq. (5). Below are tasks that we believe can be effectively represented and enhanced through this framework.
8.1.1 In-Hand Manipulation
For in-hand manipulation, our framework enables instantaneously incomplete cages to dynamically form complete cages over time, guiding objects toward target states with containment guarantees, significantly relaxing the requirement for in-hand caging (Bircher et al., 2021) in both quasistatic and dynamic cases.
In quasi-static cases, the state space is $\mathcal{S}_{obj}\subset SE(3)$, encoding object pose. The cage can form in two ways: through geometric constraints, where spatial finger contact arrangements prevent the object from escaping, or through energy-based constraints using potential energy functions where the energy boundary forms the cage as defined in Eq. (11): $\mathcal{S}_{cage}^{t}=\{\mathbf{q}_{t}\in SE(3)|E(\mathbf{q}_{t})<E_{max}\},$ where $E(\mathbf{q}_{t})$ represents the combination of gravitational potential energy and virtual potential fields that model contact constraints. These two approaches can also work in tandem -— potential fields control object motions while geometric constraints provide explicit caging boundaries, as demonstrated in (Bircher et al., 2021)
In dynamic cases, the state $\mathbf{q}_{t}$ becomes $(\mathbf{x}_{t},\dot{\mathbf{x}}_{t})$ and the state space expands to $\mathcal{S}_{obj}\subset SE(3)\times se(3)$ to incorporate both the pose and the twist, where the cage emerges as energy barriers. Similarly in Eq. (12), the energy-based cage can be explicitly defined as: $\mathcal{S}_{cage}^{t}=\{(\mathbf{x}_{t},\dot{\mathbf{x}}_{t})\in SE(3)\times se(3)|E_{p}(\mathbf{x}_{t})+E_{k}(\dot{\mathbf{x}}_{t})<E_{max}\},$ where $E_{p}(\mathbf{x}_{t})$ represents the potential energy from gravity and contact forces, and $E_{k}(\dot{\mathbf{x}}_{t})=\frac{1}{2}\dot{\mathbf{x}}_{t}^{T}\mathbf{M}\dot{\mathbf{x}}_{t}$ is the kinetic energy determined by the object’s velocity and inertia matrix $\mathbf{M}$.
Maintaining this time-varying cage requires coordinated finger actions including sliding for continuous contact adjustment, rolling for smooth surface transitions, and gaiting for discrete contact reestablishment. These motion primitives and underlying finger reconfigurations strategically evolve both geometric constraints and energy barriers over time, known as Caging in Time.
8.1.2 Extrinsic Dexterity
Extrinsic dexterity leverages environmental features as complementary cage elements, extending manipulation capabilities beyond what is possible with end-effectors alone.
As aforementioned, in quasi-static setup, the state space $\mathcal{S}_{obj}\subset SE(3)$ represents the object pose, while the cage is defined geometrically through the joint arrangement of the end-effector and environmental features like surfaces, corners, and obstacles. Similarly to Eq. (3), these elements collectively form spatial barriers around $\mathcal{C}_{free}^{obj}$, preventing uncontrolled object motion while considering the potential energy from gravity and interactions (Mahler et al., 2016).
Also in dynamic scenarios, the state space expands to $\mathcal{S}_{obj}\subset SE(3)\times se(3)$. Similarly to the energy-based cage for in-hand manipulation, the cage can be defined as $\mathcal{S}_{cage}^{t}=\{(\mathbf{q}_{t},\dot{\mathbf{q}}_{t})\in SE(3)\times se(3)|E_{p}(\mathbf{q}_{t})+E_{k}(\dot{\mathbf{q}}_{t})<E_{max}\}$, where the kinetic energy of the controlled object momentum, the potential energy of the gravitational effects and the contact forces together create energy barriers that guide the motion of the object within admissible regions, allowing manipulation with fewer contacts.
The cage is maintained through strategic management of object-environment and object-robot contacts relative to environmental features, and multi-modal motion primitives like pushing, flipping and grasping. Unlike traditional extrinsic dexterity approaches (Yang et al., 2024), Caging in Time explicitly plans transitions between contact states without requiring continuous feedback, extensive training, or specific object geometry limits, ensuring continuous caging in time despite uncertainties.
8.1.3 Deformable Object Manipulation
For deformable objects, each state $\mathbf{q}_{t}\subset\mathbb{R}^{3}$ directly represents the object’s continuous shape in $\mathbb{R}^{3}$ within the state space $\mathcal{S}_{obj}\subset 2^{\mathbb{R}^{3}}$, rather than a discrete configuration, naturally aligning with PSS representation. These shapes can be encoded through parametric approaches with predefined models (Pokorny et al., 2013), or nonparametric methods using point distributions (Shi et al., 2023). In cases where objects are considered non-elastic, they maintain their deformed shape after contact is removed, allowing representation within $\mathbb{R}^{3}$ without modeling time-dependent recovery dynamics.
The cage forms through strategically distributed contacts that constrain the PSS while transforming the object shape. Using an estimated deformation model that can cover general PSS motion similarly to Eq. (9) in planar pushing, we can create spatial barriers that limit possible deformations from exceeding the cage, enabling shape control without requiring precise physical models of complex deformation or large amounts of data.
Unlike object pushing, the action space for deformable object manipulation encompasses material-specific motion primitives such as pushing, pinching, folding, and rolling. By sequencing these primitives strategically, Caging in Time creates virtual cages whose deformation directly controls object transformation, maintaining reliable manipulation even under occlusion or complex deformation scenarios.
8.2 Implementation Guidelines
Adapting Caging in Time to new manipulation applications requires systematic consideration of several key elements:
State Space Formulation: Define $\mathcal{S}_{obj}$ to capture task-relevant properties as mentioned above, along with state propagation function discussed in Sec. 4.2:
$\mathcal{Q}_{t+1}$ $=\Pi(\mathcal{QV}_{t})$
Note that the propagation function $\Pi$ can be implemented either through analytical models as presented in this work or data-driven predictive frameworks such as neural networks and diffusion-based models.
Action Space Design: Define the action space $\mathcal{U}$ satisfying $\forall t,\exists u_{t}\in\mathcal{U}:\Pi(\mathcal{QV}_{t})\subset\mathcal{S}^{t}_{cage}$. The action space can be discrete as in our pushing tasks or continuous as in our ball balancing task with tilting angles, depending on task requirements and the control strategy.
Cage Region: Design the cage in time $\mathcal{S}^{t}_{cage},t=0,1,\dots,T$ to balance precision and robustness while ensuring the caging condition $\mathcal{Q}_{t+1}=\Pi(\mathcal{QV}_{t})\subset\mathcal{S}^{t}_{cage}$ holds. While our work mainly employed shape invariant cages, Caging in Time naturally supports time-varying cage morphology where $\mathcal{S}^{t}_{cage}$ can reshape to accommodate state-dependent constraints, enabling wider applications such as in-hand manipulation or deformable object manipulation.
Action Selection: Determining optimal actions $u_{t}\in\mathcal{U}$ requires a strategy that satisfies $\Pi(\mathcal{QV}_{t})\subset\mathcal{S}^{t}_{cage}$. Our approach employed optimization-based methods like exhaustive search for planar pushing or quadratic programming for ball manipulation. Alternative paradigms such as reinforcement learning could extend applicability to more scenarios. Importantly, as discussed in Alg. 1 and Fig. 21, hardware constraints must be incorporated to ensure the actions can be deployed in real-world physical systems.
Computational efficiency: Current action selection and PSS propagation require $26.2\pm 8.5$ ms (pushing) and $67.3\pm 12.6$ ms (ball balancing) on a single CPU thread, with potential for optimization through GPU acceleration. Though currently offline-computed, these timings show feasibility for integrating Caging in Time into the online planning for more diverse and dynamic scenarios.
@@ 9 Conclusion
In this work, we proposed and evaluated Caging in Time, a novel theory for robust object manipulation. Our framework demonstrated robust performance in both quasi-static and dynamic tasks without requiring detailed object information or real-time feedback. Rigorous evaluations highlighted the framework’s resilience and adaptability to various objects and dynamic scenarios. The Caging in Time approach proved effective in handling new objects, positional perturbations, and challenging dynamic tasks, showcasing its potential for reliable manipulation in uncertain environments.
@@
Looking forward, we aim to address these limitations and expand the horizons of Caging in Time. A key direction for future research is the integration of learning techniques, LLMs, and diffusion models to autonomously construct and learn Caging in Time tasks. This approach could enable the framework to automatically derive appropriate propagation functions and strategies, reducing the need for manual parameter tuning. We also plan to explore its applications as mentioned in Sec. 8, while potentially leveraging reinforcement learning to handle increased task complexity and environmental variability.
@@ 9.1.1 The Peshkin Distance
where $a$ represents the radius of the object’s bounding circle centered at the mass center, $c$ is the distance from the mass center to the contact point, and $\beta$ is the angle subtended by the motion direction and the line connecting the contact point with the mass center. The angles $\beta_{0}$ and $\beta_{1}$ correspond to the initial and final orientation angles for this rotation.
@@
$=\frac{mc^{2}}{a^{2}+c^{2}}\sin{\beta_{0}}\cos{\beta_{0}}\leq m\sin{\beta_{0}}\cos{\beta_{0}}\leq\frac{m}{2}$
@@
$\mathcal{V}_{q_{t}}=U(q_{t},u_{t})=\text{if $\mathbf{d}(q_{t},\mathcal{L})>r+d_{push}$ or $u_{t}$ is {None}:}$ (33)
@@ 9.2.1 System Dynamics
The system dynamics of the ball on the plate can be described by the following equation:
@@
$=\frac{m(1+\eta_{m})}{m+\frac{I_{b}}{r_{b}^{2}}}\Big{(}(I_{n+1}-\mathbf{n}(\boldsymbol{\theta})\mathbf{n}(\boldsymbol{\theta})^{T})$
×$\quad\times(\mathbf{g}_{\theta}+\mathbf{a}_{p}+\boldsymbol{\eta}_{p})\Big{)}-(\mu_{r}+\eta_{\mu})\dot{\mathbf{x}}_{t}$
@@
$f_{t}=$ $\frac{m}{m+\frac{I_{b}}{r_{b}^{2}}}(1+\eta_{m})((g+(\ddot{z}_{p}+\eta_{z}))\sin\theta+$ (38)
$(\ddot{x}_{p}+\eta_{x})\cos\theta)-(\mu_{r}+\eta_{\mu})\dot{x}_{t}$
$g_{t}=$ $\frac{m}{m+\frac{I_{b}}{r_{b}^{2}}}(1+\eta_{m})((g+(\ddot{z}_{p}+\eta_{z}))\cos\theta-$
$(\ddot{x}_{p}+\eta_{x})\sin\theta)$